I need help solving sytems of equations word problems. An example of one is: Sally makes dolls for a living. She sells dolls for 21$ each. Her expenses for producing the dolls are 1000$ for the sewing machine plus 12$ per doll for materials. How many dolls must she sell to break even?
Answer by rfer(16322)

12x+1000=21x
12x-21x=-1000
-9x=-1000
x=-1000/-9
x=111.1=112 dolls to break even